A trench mortar fires a projectile at an angle of 53° above the horizontal with a muzzle velocity of 60 m/s. A tank is advancing directly towards the mortar on level ground at a speed of 3 m/s. What should be the distance from the mortar to tank at the instant the mortar is fired in order to score a hit?


1
Expert's answer
2021-06-18T07:54:45-0400

R = "\dfrac{u²\sin2\theta}{g}=\dfrac{60²\sin106°}{10}= 346m"


"T = \dfrac{2u\sin\theta}{g} = \dfrac{2×60\sin53°}{10}= 9.6s"


"s= ut = 3\u00d7 9.6 = 28.8m"


"\therefore" The distance from the mortar to tank at the instant the mortar is fired is (28.8m + 9.6m = ) 38.4m.